\begin{thebibliography}{25}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Ben{-}David et~al.(2010{\natexlab{a}})Ben{-}David, Blitzer, Crammer,
  Kulesza, Pereira, and Vaughan]{DavidBCKPV10}
Ben{-}David, S., Blitzer, J., Crammer, K., Kulesza, A., Pereira, F., and
  Vaughan, J.~W.
\newblock A theory of learning from different domains.
\newblock \emph{Mach. Learn.}, 79\penalty0 (1-2):\penalty0 151--175,
  2010{\natexlab{a}}.
\newblock \doi{10.1007/s10994-009-5152-4}.
\newblock URL \url{https://doi.org/10.1007/s10994-009-5152-4}.

\bibitem[Ben{-}David et~al.(2010{\natexlab{b}})Ben{-}David, Lu, Luu, and
  P{\'{a}}l]{DavidLLP10}
Ben{-}David, S., Lu, T., Luu, T., and P{\'{a}}l, D.
\newblock Impossibility theorems for domain adaptation.
\newblock In Teh, Y.~W. and Titterington, D.~M. (eds.), \emph{Proceedings of
  the Thirteenth International Conference on Artificial Intelligence and
  Statistics, {AISTATS} 2010, Chia Laguna Resort, Sardinia, Italy, May 13-15,
  2010}, volume~9 of \emph{{JMLR} Proceedings}, pp.\  129--136. JMLR.org,
  2010{\natexlab{b}}.
\newblock URL \url{http://proceedings.mlr.press/v9/david10a.html}.

\bibitem[Ganin \& Lempitsky(2015)Ganin and Lempitsky]{GaninL15}
Ganin, Y. and Lempitsky, V.~S.
\newblock Unsupervised domain adaptation by backpropagation.
\newblock In Bach, F.~R. and Blei, D.~M. (eds.), \emph{Proceedings of the 32nd
  International Conference on Machine Learning, {ICML} 2015, Lille, France,
  6-11 July 2015}, volume~37 of \emph{{JMLR} Workshop and Conference
  Proceedings}, pp.\  1180--1189. JMLR.org, 2015.
\newblock URL \url{http://proceedings.mlr.press/v37/ganin15.html}.

\bibitem[Germain et~al.(2013)Germain, Habrard, Laviolette, and
  Morvant]{GermainHLM13}
Germain, P., Habrard, A., Laviolette, F., and Morvant, E.
\newblock A pac-bayesian approach for domain adaptation with specialization to
  linear classifiers.
\newblock In \emph{Proceedings of the 30th International Conference on Machine
  Learning, {ICML} 2013, Atlanta, GA, USA, 16-21 June 2013}, volume~28 of
  \emph{{JMLR} Workshop and Conference Proceedings}, pp.\  738--746. JMLR.org,
  2013.
\newblock URL \url{http://proceedings.mlr.press/v28/germain13.html}.

\bibitem[Goodfellow et~al.(2015)Goodfellow, Shlens, and
  Szegedy]{GoodfellowSS14}
Goodfellow, I.~J., Shlens, J., and Szegedy, C.
\newblock Explaining and harnessing adversarial examples.
\newblock In Bengio, Y. and LeCun, Y. (eds.), \emph{3rd International
  Conference on Learning Representations, {ICLR} 2015, San Diego, CA, USA, May
  7-9, 2015, Conference Track Proceedings}, 2015.
\newblock URL \url{http://arxiv.org/abs/1412.6572}.

\bibitem[Gretton et~al.(2012)Gretton, Borgwardt, Rasch, Sch{\"{o}}lkopf, and
  Smola]{mmd}
Gretton, A., Borgwardt, K.~M., Rasch, M.~J., Sch{\"{o}}lkopf, B., and Smola,
  A.~J.
\newblock A kernel two-sample test.
\newblock \emph{J. Mach. Learn. Res.}, 13:\penalty0 723--773, 2012.
\newblock URL \url{http://dl.acm.org/citation.cfm?id=2188410}.

\bibitem[Guo et~al.(2020)Guo, Li, and Chen]{linearization}
Guo, Y., Li, Q., and Chen, H.
\newblock Backpropagating linearly improves transferability of adversarial
  examples.
\newblock In Larochelle, H., Ranzato, M., Hadsell, R., Balcan, M., and Lin, H.
  (eds.), \emph{Advances in Neural Information Processing Systems 33: Annual
  Conference on Neural Information Processing Systems 2020, NeurIPS 2020,
  December 6-12, 2020, virtual}, 2020.
\newblock URL
  \url{https://proceedings.neurips.cc/paper/2020/hash/00e26af6ac3b1c1c49d7c3d79c60d000-Abstract.html}.

\bibitem[{Hull}(1994)]{uspsdataset}
{Hull}, J.~J.
\newblock A database for handwritten text recognition research.
\newblock \emph{IEEE Transactions on Pattern Analysis and Machine
  Intelligence}, 16\penalty0 (5):\penalty0 550--554, 1994.
\newblock \doi{10.1109/34.291440}.

\bibitem[Kifer et~al.(2004)Kifer, Ben{-}David, and Gehrke]{KiferBG04}
Kifer, D., Ben{-}David, S., and Gehrke, J.
\newblock Detecting change in data streams.
\newblock In Nascimento, M.~A., {\"{O}}zsu, M.~T., Kossmann, D., Miller, R.~J.,
  Blakeley, J.~A., and Schiefer, K.~B. (eds.), \emph{(e)Proceedings of the
  Thirtieth International Conference on Very Large Data Bases, {VLDB} 2004,
  Toronto, Canada, August 31 - September 3 2004}, pp.\  180--191. Morgan
  Kaufmann, 2004.
\newblock \doi{10.1016/B978-012088469-8.50019-X}.
\newblock URL \url{http://www.vldb.org/conf/2004/RS5P1.PDF}.

\bibitem[Kingma \& Ba(2015)Kingma and Ba]{KingmaB14}
Kingma, D.~P. and Ba, J.
\newblock Adam: {A} method for stochastic optimization.
\newblock In Bengio, Y. and LeCun, Y. (eds.), \emph{3rd International
  Conference on Learning Representations, {ICLR} 2015, San Diego, CA, USA, May
  7-9, 2015, Conference Track Proceedings}, 2015.
\newblock URL \url{http://arxiv.org/abs/1412.6980}.

\bibitem[LeCun \& Cortes(2010)LeCun and Cortes]{mnisthandwrittendigit}
LeCun, Y. and Cortes, C.
\newblock {MNIST} handwritten digit database.
\newblock 2010.
\newblock URL \url{http://yann.lecun.com/exdb/mnist/}.

\bibitem[Liang et~al.(2020)Liang, Hu, and Feng]{SHOT}
Liang, J., Hu, D., and Feng, J.
\newblock Do we really need to access the source data? source hypothesis
  transfer for unsupervised domain adaptation.
\newblock In \emph{Proceedings of the 37th International Conference on Machine
  Learning, {ICML} 2020, 13-18 July 2020, Virtual Event}, volume 119 of
  \emph{Proceedings of Machine Learning Research}, pp.\  6028--6039. {PMLR},
  2020.
\newblock URL \url{http://proceedings.mlr.press/v119/liang20a.html}.

\bibitem[Long et~al.(2015)Long, Cao, Wang, and Jordan]{DAN}
Long, M., Cao, Y., Wang, J., and Jordan, M.~I.
\newblock Learning transferable features with deep adaptation networks.
\newblock In Bach, F.~R. and Blei, D.~M. (eds.), \emph{Proceedings of the 32nd
  International Conference on Machine Learning, {ICML} 2015, Lille, France,
  6-11 July 2015}, volume~37 of \emph{{JMLR} Workshop and Conference
  Proceedings}, pp.\  97--105. JMLR.org, 2015.
\newblock URL \url{http://proceedings.mlr.press/v37/long15.html}.

\bibitem[Long et~al.(2017)Long, Zhu, Wang, and Jordan]{JAN}
Long, M., Zhu, H., Wang, J., and Jordan, M.~I.
\newblock Deep transfer learning with joint adaptation networks.
\newblock In Precup, D. and Teh, Y.~W. (eds.), \emph{Proceedings of the 34th
  International Conference on Machine Learning, {ICML} 2017, Sydney, NSW,
  Australia, 6-11 August 2017}, volume~70 of \emph{Proceedings of Machine
  Learning Research}, pp.\  2208--2217. {PMLR}, 2017.
\newblock URL \url{http://proceedings.mlr.press/v70/long17a.html}.

\bibitem[Long et~al.(2018)Long, Cao, Wang, and Jordan]{cdan}
Long, M., Cao, Z., Wang, J., and Jordan, M.~I.
\newblock Conditional adversarial domain adaptation.
\newblock In Bengio, S., Wallach, H.~M., Larochelle, H., Grauman, K.,
  Cesa{-}Bianchi, N., and Garnett, R. (eds.), \emph{Advances in Neural
  Information Processing Systems 31: Annual Conference on Neural Information
  Processing Systems 2018, NeurIPS 2018, December 3-8, 2018, Montr{\'{e}}al,
  Canada}, pp.\  1647--1657, 2018.
\newblock URL
  \url{https://proceedings.neurips.cc/paper/2018/hash/ab88b15733f543179858600245108dd8-Abstract.html}.

\bibitem[Luo et~al.(2018)Luo, Liu, Wei, and Xu]{LuoLWX18}
Luo, B., Liu, Y., Wei, L., and Xu, Q.
\newblock Towards imperceptible and robust adversarial example attacks against
  neural networks.
\newblock In McIlraith, S.~A. and Weinberger, K.~Q. (eds.), \emph{Proceedings
  of the Thirty-Second {AAAI} Conference on Artificial Intelligence, (AAAI-18),
  the 30th innovative Applications of Artificial Intelligence (IAAI-18), and
  the 8th {AAAI} Symposium on Educational Advances in Artificial Intelligence
  (EAAI-18), New Orleans, Louisiana, USA, February 2-7, 2018}, pp.\
  1652--1659. {AAAI} Press, 2018.
\newblock URL
  \url{https://www.aaai.org/ocs/index.php/AAAI/AAAI18/paper/view/16217}.

\bibitem[Madry et~al.(2018)Madry, Makelov, Schmidt, Tsipras, and
  Vladu]{MadryMSTV18}
Madry, A., Makelov, A., Schmidt, L., Tsipras, D., and Vladu, A.
\newblock Towards deep learning models resistant to adversarial attacks.
\newblock In \emph{6th International Conference on Learning Representations,
  {ICLR} 2018, Vancouver, BC, Canada, April 30 - May 3, 2018, Conference Track
  Proceedings}, 2018.
\newblock URL \url{https://openreview.net/forum?id=rJzIBfZAb}.

\bibitem[Moosavi{-}Dezfooli et~al.(2016)Moosavi{-}Dezfooli, Fawzi, and
  Frossard]{Moosavi-Dezfooli16}
Moosavi{-}Dezfooli, S., Fawzi, A., and Frossard, P.
\newblock Deepfool: {A} simple and accurate method to fool deep neural
  networks.
\newblock In \emph{2016 {IEEE} Conference on Computer Vision and Pattern
  Recognition, {CVPR} 2016, Las Vegas, NV, USA, June 27-30, 2016}, pp.\
  2574--2582. {IEEE} Computer Society, 2016.
\newblock \doi{10.1109/CVPR.2016.282}.
\newblock URL \url{https://doi.org/10.1109/CVPR.2016.282}.

\bibitem[Papernot et~al.(2016)Papernot, McDaniel, and
  Goodfellow]{DBLP:journals/corr/PapernotMG16}
Papernot, N., McDaniel, P.~D., and Goodfellow, I.~J.
\newblock Transferability in machine learning: from phenomena to black-box
  attacks using adversarial samples.
\newblock \emph{CoRR}, abs/1605.07277, 2016.
\newblock URL \url{http://arxiv.org/abs/1605.07277}.

\bibitem[Petrov \& Hospedales(2019)Petrov and Hospedales]{Transfer}
Petrov, D. and Hospedales, T.~M.
\newblock Measuring the transferability of adversarial examples.
\newblock \emph{CoRR}, abs/1907.06291, 2019.
\newblock URL \url{http://arxiv.org/abs/1907.06291}.

\bibitem[Szegedy et~al.(2014)Szegedy, Zaremba, Sutskever, Bruna, Erhan,
  Goodfellow, and Fergus]{DBLP:journals/corr/SzegedyZSBEGF13}
Szegedy, C., Zaremba, W., Sutskever, I., Bruna, J., Erhan, D., Goodfellow,
  I.~J., and Fergus, R.
\newblock Intriguing properties of neural networks.
\newblock In Bengio, Y. and LeCun, Y. (eds.), \emph{2nd International
  Conference on Learning Representations, {ICLR} 2014, Banff, AB, Canada, April
  14-16, 2014, Conference Track Proceedings}, 2014.
\newblock URL \url{http://arxiv.org/abs/1312.6199}.

\bibitem[Tousch \& Renaudin(2020)Tousch and Renaudin]{adalib2020}
Tousch, A.-M. and Renaudin, C.
\newblock (yet) another domain adaptation library, 2020.
\newblock URL \url{https://github.com/criteo-research/pytorch-ada}.

\bibitem[Venkateswara et~al.(2017{\natexlab{a}})Venkateswara, Eusebio,
  Chakraborty, and Panchanathan]{officehome}
Venkateswara, H., Eusebio, J., Chakraborty, S., and Panchanathan, S.
\newblock Deep hashing network for unsupervised domain adaptation.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pp.\  5018--5027, 2017{\natexlab{a}}.

\bibitem[Venkateswara et~al.(2017{\natexlab{b}})Venkateswara, Eusebio,
  Chakraborty, and Panchanathan]{venkateswara2017deep}
Venkateswara, H., Eusebio, J., Chakraborty, S., and Panchanathan, S.
\newblock Deep hashing network for unsupervised domain adaptation.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition}, pp.\  5018--5027, 2017{\natexlab{b}}.

\bibitem[Wang \& Deng(2018)Wang and Deng]{DAsurvey}
Wang, M. and Deng, W.
\newblock Deep visual domain adaptation: {A} survey.
\newblock \emph{Neurocomputing}, 312:\penalty0 135--153, 2018.
\newblock \doi{10.1016/j.neucom.2018.05.083}.
\newblock URL \url{https://doi.org/10.1016/j.neucom.2018.05.083}.

\end{thebibliography}

\documentclass[nohyperref]{article}

\usepackage{microtype}
\usepackage{graphicx}
\usepackage{booktabs} 

\usepackage{hyperref}

 \usepackage[accepted]{icml2022}

\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{mathtools}
\usepackage{amsthm}
\usepackage{subfig}

\usepackage[capitalize,noabbrev]{cleveref}

\theoremstyle{plain}

\theoremstyle{definition}

\theoremstyle{remark}

\usepackage[textsize=tiny]{todonotes}

\icmltitlerunning{Multi-step domain adaptation by adversarial attack to $\mathcal{H} \Delta \mathcal{H}$-divergence}

\begin{document}

\twocolumn[
\icmltitle{Multi-step domain adaptation by adversarial attack to $\mathcal{H} \Delta \mathcal{H}$-divergence}



\begin{icmlauthorlist}
\icmlauthor{Arip Asadulaev}{Itmo,Airi}
\icmlauthor{Alexander Panfilov}{Tubingen}
\icmlauthor{Andrey Filchenkov}{Itmo}
\end{icmlauthorlist}

\icmlaffiliation{Itmo}{ITMO University, Saint-Petersburg, Russia}
\icmlaffiliation{Airi}{Artificial Intelligence Research Institute, Moscow, Russia}
\icmlaffiliation{Tubingen}{Tubingen Universitat, Germany}

\icmlcorrespondingauthor{Arip Asadulaev}{aripasadulaev@itmo.ru}

\icmlkeywords{Machine Learning, ICML}

\vskip 0.3in
]



\printAffiliationsAndNotice{} 

\begin{abstract}
Adversarial examples are transferable between different models. In our paper, we propose to use this property for multi-step domain adaptation. In unsupervised domain adaptation settings, we demonstrate that replacing the source domain with adversarial examples to $\mathcal{H} \Delta \mathcal{H}$-divergence can improve source classifier accuracy on the target domain. Our method can be connected to most domain adaptation techniques. We conducted a range of experiments and achieved improvement in accuracy on Digits and Office-Home datasets. 
\end{abstract}\vspace{-4mm}
\vspace{-5mm}\section{Introduction}
In domain adaptation, we train a model to make correct predictions on the different domains named source and target~\cite{DavidBCKPV10}. Domain adaptation techniques aim to bring the target domain ``closer‘’ to the source domain or find a common representation for these domains.~\cite{GaninL15, cdan}. 

Theoretically, the accuracy of domain adaptation is bounded by the source domain error and the divergence $\mathcal{H} \Delta \mathcal{H}$ between source and target domains~\cite{DavidBCKPV10,DavidLLP10,GermainHLM13}.  The divergence $\mathcal{H} \Delta \mathcal{H}$ can be approximated by the binary classifier between source and target domains. We find that source class samples turned into the target class sample (by an adversarial attack to the $\mathcal{H} \Delta \mathcal{H}$ neural approximation) are actually "closer" to the real target samples. We show that training the source classifier on these samples can improve its accuracy in the target domain. 

We call our method $\mathcal{H} \Delta \mathcal{H}$-divergence Domain Adaptation (HDA). We tested a range of domain adaptation algorithms coupled with the proposed technique on datasets ``Digits’’~\cite{mnisthandwrittendigit} and ``Office-Home’’~\cite{officehome}. 

\section{Background}
\textbf{Adversarial Attacks:} Having  sample $x$, target label $y$, model parameters $\theta$, loss function $L$, we can apply adversarial attack~\cite{DBLP:journals/corr/SzegedyZSBEGF13,GoodfellowSS14,Moosavi-Dezfooli16,MadryMSTV18,LuoLWX18}. For example, Fast Sign Gradient Descent (FSGD)~\cite{GoodfellowSS14} can be presented  as: ${x}_{i+1}={x}_{i}+\varepsilon \operatorname{sign}\left(-\nabla_{x} L\left(\theta, {x}_{i}, y_{i}\right)\right)$, with perturbation size $\varepsilon$ around the original image $x_i$. It was shown that  adversarial examples are transferable between different models~\cite{Transfer}. Recent studies have revealed that the adversarial examples are largely invariant to the models trained on different sampled datasets~\cite{DBLP:journals/corr/PapernotMG16}. 

\begin{table*}[h!]
\centering
\vskip 0.1in
\begin{center}
\begin{footnotesize}
\begin{sc}

\begin{tabular}{lllllll} 
\hline

Model       & MNIST$\rightarrow$MNIST-M & MNIST-M$\rightarrow$MNIST      & SVHN$\rightarrow$MNIST         & MNIST$\rightarrow$SVHN        &USPS$\rightarrow$MNIST \\ \hline

DANN        & 97.2 +- 1.37          & \textbf{73.4 +- 1.71}                  & 21.3 +- 2.38                  & 67.3 +- 3.20                           & 97.2 +- 1.37            \\ 
CDAN        & 97.8 +- 0.22          & 61.0 +- 2.76                    & 15.7 +- 4.55                           & 61.5 +- 5.14                          & 97.8 +- 0.22            \\ 
CDAN-E      & 97.8 +- 0.41          & 67.0 +- 4.32                    & 13.9 +- 1.47                           & 60.5 +- 4.44                         & 97.8 +- 0.41             \\ 
DAN         & 97.6 +- 1.17          & 43.6 +- 1.93                    & 19.2 +- 2.97                           & 61.2 +- 2.27                         & 97.6 +- 1.17           \\ 
JAN         & 97.6 +- 0.39          & 44.7 +- 4.48                   & 10.7 +- 1.70                            & 59.9 +- 3.14                     & 97.6 +- 0.39         \\ 
SHOT        & 98.1 +- 1.22          & 77.4 +- 1.44                   & 24.4 +- 1.50                          & 98.9 +- 2.11                         &98.0 +- 0.51           \\ \hline

DANN + HDA  & \textbf{98.1 +- 0.25} & 70.7 +- 2.86                  & \textbf{24.5 +- 3.47}                  & \textbf{78.6 +- 7.99}                  & \textbf{98.1 +- 0.25}            \\ 
CDAN + HDA  & \textbf{98.0 +- 0.41} & \textbf{65.2 +- 3.69}         & \textbf{17.2 +- 2.24}                  & \textbf{70.8 +- 8.29}                 & \textbf{98.0 +- 0.41}            \\ 
CDAN-E + HDA & \textbf{98.2 +- 0.03} & \textbf{69.9 +- 2.28}         & \textbf{16.0 +- 1.46}                  & \textbf{72.8 +- 3.95}                  & \textbf{98.2 +- 0.03}           \\ 
DAN + HDA   & \textbf{97.6 +- 0.15}  & \textbf{44.6 +- 1.24}         & \textbf{23.2 +- 2.91}                  & 60.7 +- 3.34                  & 97.6 +- 0.15            \\ 
JAN + HDA   & 97.1 +- 0.16          & \textbf{49.8 +- 6.47}         & \textbf{13.6 +- 2.18}                  & 59.3 +- 10.56                  & 97.1 +- 0.16            \\ 
SHOT + HDA  & \textbf{98.2 +- 0.45}          &  \textbf{77.6 +- 1.91}                            & \textbf{42.6 +- 3.21}                  & 98.9 +- 1.72                & 98.0 +- 0.40            \\ \hline
\end{tabular}
\end{sc}
\end{footnotesize}
\end{center}

\caption{Results of domain adaptation with HDA on Digits datasets.}
\label{tab:table-1}
\end{table*}
\begin{table*}[h!]
\centering
\begin{center}
\begin{footnotesize}
\begin{sc}
\begin{tabular}{llllllllllllll}
\hline
Model           & A$\rightarrow$C & A$\rightarrow$P & A$\rightarrow$R & C$\rightarrow$A & C$\rightarrow$P & C$\rightarrow$R 
                & P$\rightarrow$A & P$\rightarrow$C & P$\rightarrow$R & R$\rightarrow$A & R$\rightarrow$C & R$\rightarrow$P \\ \hline
DANN            & 45.6 & 59.3 & 70.1 & 47.0 & 58.5 & 60.9 & 46.1 & 43.7 & 68.5 & 63.2 & 51.8 & 76.8 \\ 
DAN            & 43.6 & 57.0 & 67.9 & 45.8 & 56.5 & 60.4 & 44.0 & 43.6 & 67.7 & 63.1 & 51.5 & 74.3\\ 
CDAN-E           & 50.7 & 70.6 & 76.0 & 57.6 & 70.0 & 70.0 & 57.4 & 50.9 & 77.3 & 70.9 & 56.7 & 81.6  \\ 
SHOT            & 57.1 & 78.1 & 81.5 & 68.0 & 78.2 & 78.1 & 67.4 & 54.9 & 82.2 & 73.3 & 58.8 & 84.3 \\ \hline
SHOT+HDA       & 57.0 & \textbf{78.5} & 80.6 & \textbf{69.1} & \textbf{79.5} & \textbf{78.4} &\textbf{68.3} & \textbf{55.5} & 81.5 & \textbf{73.7} & \textbf{60.5} & 83.8 \\ \hline
\end{tabular}
\end{sc}
\end{footnotesize}
\end{center}
\caption{Results on Office-Home datasets. Domains are: Art (A), Clipart (C), Product (P), Real-World (R)}
\label{tab:table-2}
\end{table*}

\textbf{Domain Adaptation:} For the given two domains, source $S_X$ and target $T_X$ over the input space $X$, the goal is to bring the target domain or its representation ``closer‘’ to the source domain. Following the learning theory~\cite{KiferBG04, DavidBCKPV10,DavidLLP10}, the domain adaptation error is bounded by the source domain classifier error and the $\mathcal{H} \Delta \mathcal{H}$ discrepancy between source and target domains. It is impossible to estimate exactly the value of $\mathcal{H} \Delta \mathcal{H}$, but it can be approximated by a binary classifier that discriminates the source and the target samples~\cite{DavidBCKPV10,DavidLLP10}. 

\section{Algorithm}

First of all, we train a binary classifier $\mathcal{H}_\omega$ between domains $S_X$ and $T_X$ that approximates $\mathcal{H} \Delta \mathcal{H}$. Then, we apply perturbations in the input space of $\mathcal{H}_\omega$ to make source samples $S_X$ look like the target domain samples $T_X$ for $\mathcal{H}_\omega$ and get a new domain $A_X$. In the third step, we pretrain the source classifier $\mathcal{F}_\theta$ on the resulted domain $A_X$, instead of the real source $S_X$. Due to the adversarial attacks' transferability, $A_X$ adversarial examples are transferable to the $\mathcal{F}_\theta$ classifier too, and actually, increase $\mathcal{F}_\theta$ accuracy on the target domain. Finally, we apply domain adaptation for the source classifier $\mathcal{F}_\theta$, using the arbitrary distance-based domain adaptation techniques~\cite{GaninL15,cdan, DAN, JAN}, but also replacing the real source data $S_X$ with the $A_X$ in the training pipeline. As a result, we obtain a multi-step domain adaptation pipeline~\cite{DAsurvey}, where at the first step, the algorithm creates a new dataset that is used by the second step method. 

\section{Experiments}
\textbf{Datasets:} All experiments were conducted in unsupervised settings (only source labels are known) on the range of digits domains (MNIST~\cite{mnisthandwrittendigit}, USPS ~\cite{uspsdataset}, MNIST-M, SVHN) and  Office-Home~\cite{venkateswara2017deep} domains.

\textbf{Settings:} In our experiments, we used $A_X$ instead of the real source $S_X$ for prominent adversarial-based approach's DANN~\cite{GaninL15}, CDAN, CDAN-E~\cite{cdan}. Also, we tested our method on Maximum Mean Discrepancy (MMD)~\cite{mmd} based on domain adaptation techniques like DAN~\cite{DAN} and JAN~\cite{JAN}. In addition, we tested our method in connection to the SHOT~\cite{SHOT} method on the Office-Home domains. We used ADA~\cite{adalib2020} library for training and testing, and all training settings were set equal to the default parameters proposed by ADA.  

The $\mathcal{H}_\omega$ domains classifier was trained 5 epochs using Adam optimizer \cite{KingmaB14} with a $0.01$ learning rate.  For $A_X$ dataset generation $l_inf$ FSGD attack with perturbation size $\varepsilon$ equal to $0.01$ and 7 steps was used. 

Before applying domain adaptation, we pretrained the $\mathcal{F}_\theta$ classifier by 10 epochs on the $A_X$ dataset. Then, each domain adaptation method used 20 epochs of training using the target domain samples. For parameters updating during the domain adaptation, we used the Adam optimizer with a learning rate equal to $0.01$. All hyperparameters were equal for each tested domain adaptation method and each pair of source and target domains.

\textbf{Results:} Averaged results over three random seeds for the best-selected hyperparameters for each adaptation task are presented in Table~\ref{tab:table-1},\ref{tab:table-2}. Across all benchmarks, the training with a $A_x$-adversarial domain makes reasonable gains over the basic settings. We find that different domains require different sizes of perturbations to generate transferable examples. Searching adversarial attack parameters more carefully individually for each task can increase adaptation accuracy.

\section{Conclusion and future works}
Our paper is still a work in progress, but we propose a simple remedy to improve the accuracy of domain adaptation methods. The proposed method allows minimizing the $\mathcal{H} \Delta \mathcal{H}$-distance between the source and target domains more successfully. By combining our method with domain adaptation techniques, we hope it may probably result in the development of less complicated and more efficient domain adaptation techniques. 
In the future, we plan to test our method on other models and other datasets, and also we are going to increase the $\mathcal{H} \Delta \mathcal{H}$-divergence attack features transferability via linearization of backpropagation~\cite{linearization} and adversarial robustness training for robust features transfer. In addition, we are planning to test our method with more advanced techniques for adversarial attacks.


\newpage

\bibliography{example_paper}
\bibliographystyle{icml2022}

\end{document}